Diffusion Morning report 10/18/06

Ted Berg tberg at stanford.edu
Wed Oct 18 06:30:37 PDT 2006


Good Morning All,
    Looking OK this morning. Epi is still being a pain-we are looking 
into it. LTO is back up after quartz replacement. Low O2 MFC was changed 
on tube2
AMAT Epi is cycling wafers.  We may turn gases on today.  No other 
issues at this time.

Diffusion Mike,Ray, Ted 	  	  	  	  	 
  	Date: 	10/18/06 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	Comments 	  	  	 
Tylan 1 	Up 	  	  	  	 
Tylan 2 	Up 	  	  	  	 
Tylan 3 	Up 	  	  	  	 
Tylan 4 	Up 	  	  	  	 
Tylan 5-BBr3 	Up 	  	  	  	 
Tylan 6-POCl3 	Up 	  	  	  	 
Tylan 7-TEOS2 	Up 	  	8.41/30 	  	 
Tylan 8-SiGe 	Up 	  	6.72/30 	  	 
Tylan 9-Poly 	Up 	  	8.88/30 	  	 
Tylan 10-Nitride 	Up 	  	10.1/30 	  	 
Tylan 11-BPSG-LTO 	up 	  	.22/30 	  	 
Tylan 12-FGA 	Up 	  	  	  	 
  	  	  	  	  	 
Thermco 1 	Up 	  	  	  	 
Thermco 2 	Up 	  	  	  	 
Thermco Poly 	Up 	  	  	  	 
Thermco Nitride 	Up 	  	  	  	 
Epi 	down 	leak bad base 	  	  	 
  	
	
	
	
	
Burn Box 	up 	  	  	  	 
ZYGO 	Up 	  	  	  	 
  	up 	  	   	  	 
GAAS 22 	  	  	  	  	 
GAAS 23 	UP 	  	  	  	 
SiH4 for furnaces 	up 	  	  	  	 

	  	B on line 200 psi 	 changeover working well  	  	 




-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20061018/12dc2671/attachment.html>


More information about the diffusion mailing list